Please help me solve the area of a triangle. base 2 1/3 yards and height 3/4 yard. (the area of the triangle equals half the product of its base and height) I got an answer of 7/8 yards. I got it wrong.
Answer by fcabanski(1391)

Area = 1/2 * base * height.


The base is 2 1/3 yard = 7/3 yard.


1/2 * 7/3 * 3/4 (multiply top x top and bottom x bottom) = 21/24 = 7/8 square yards.


The only part you had wrong was the units.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)